Find an equation of the parabola that has vertex (1,2) and focus (-1,2).
I dont know how to solve this problem, if anybody could help me out or give me some examples or refer me to other web pages where I can find examples I would really appreciate it. Thank you.

Answer by scott8148(6628)

from wikipedia

"In Cartesian coordinates, a parabola with an axis parallel to the y axis with vertex (h, k), focus (h, k + p), and directrix y = k - p, with p being the distance from the vertex to the focus, has the equation with axis parallel to the y-axis"

(x-h)^2=4p(y-k)

p=-2, h=1, k=2 ___ (x-1)^2=-8(y-2)